OpenStudy (anonymous):

Can someone help me with another Algebra 1 question for a fan and medal?

OpenStudy (anonymous):

i can

OpenStudy (anonymous):

Ok, thank you! Hang on, give me a sec so i can get the question on here

OpenStudy (anonymous):

ok

OpenStudy (one098):

sure hannah!

OpenStudy (anonymous):

Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x).

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Im not sure why all the question marks are there, so just ignore them please

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

10th grade but taking algebra 1 online

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what was the problem again

OpenStudy (anonymous):

wait you need to ask your teacher about the ??????????????? marks

OpenStudy (anonymous):

Complete the table below to solve the equation 2.5x 10.5 = 64(0.5x)

OpenStudy (anonymous):

you should do 10.5-64

OpenStudy (anonymous):

yeah just ignore the question marks. I dont know why they showed up, but there not in the actual question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on ok

OpenStudy (anonymous):

Oh, yeah there is a - in between the 2.5x and the 10.5 so let me retype it for you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2.5x - 10.5 = 64(0.5x)

OpenStudy (anonymous):

Sorry for the confusion

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

=-26.25 +32= 32-26.25 = 5.75

OpenStudy (anonymous):

you still there

OpenStudy (anonymous):

?????????????????

OpenStudy (anonymous):

you awsner is =-26.25 +32= 32-26.25 = 5.75 ok

OpenStudy (anonymous):

Yeah sry

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

but I wouldnt i need multiple answers since there is a table?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

sry, my typing skills are not on point today... lol. but wouldnt i need multiple answers since there is a table?*

OpenStudy (anonymous):

ok hold on ok

OpenStudy (anonymous):

lol you still there

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

can you write the problem again plesae

OpenStudy (anonymous):

and with need m\times too plesae

OpenStudy (anonymous):

just hold on ok

OpenStudy (anonymous):

Hey, my food arrived. Im gonna go eat, will be back in around 10-15 minutes

OpenStudy (anonymous):

ok your anwser will be on your screen ok

OpenStudy (anonymous):

ok, thank you for your time!

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

Ok back, did you get the answer?

OpenStudy (anonymous):

yes half 32x-2.5x -10.5=29.5x

OpenStudy (anonymous):

there is more ok

OpenStudy (anonymous):

thats some

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got it i write it down ok

OpenStudy (anonymous):

can you give me some deatils about you work plesae

OpenStudy (anonymous):

Not really much details to give. Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x) and the table is:

OpenStudy (anonymous):

Ugh this question mark bs is really getting annoying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is this your homework

OpenStudy (anonymous):

Yeah, just ignore the question marks and think its a - sign lol

OpenStudy (anonymous):

yes, this is homework

OpenStudy (anonymous):

ok lol

OpenStudy (anonymous):

when is it due

OpenStudy (anonymous):

hey when is it due

OpenStudy (anonymous):

i need to known

OpenStudy (anonymous):

sry, its due tomorrow

OpenStudy (anonymous):

AHHHHHH WHATTTTTTTTTTTT

OpenStudy (anonymous):

yea, i know :/

OpenStudy (anonymous):

WHY

OpenStudy (anonymous):

i dont no, i reacted the same way in class when my teacher told us its due tomorrow

OpenStudy (anonymous):

THATS STUPID

OpenStudy (anonymous):

you need to do onlnie school easy

OpenStudy (anonymous):

THATS SOOO NOT FAIR

OpenStudy (anonymous):

i know, it really isnt :/ but this is the 2nd to last problem, so im almost done

OpenStudy (anonymous):

-clasp

OpenStudy (anonymous):

ok 2 is =2.25-10.5=-5.5 ok

OpenStudy (anonymous):

IRS 5.5

OpenStudy (anonymous):

I MEAN -5.5

OpenStudy (anonymous):

IRS?

OpenStudy (anonymous):

ITS -5.5 FOR 2 OK

OpenStudy (anonymous):

beside that is 16 ok

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

hey you want to known something

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

I AM IN 7 GRADE

OpenStudy (anonymous):

lol ik i am really smart i ddo 10 grade stuff

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

yeape

OpenStudy (anonymous):

wow, your gonna be real successful in life :O

OpenStudy (anonymous):

lol idk about that

OpenStudy (anonymous):

next one is 3.0

OpenStudy (anonymous):

wait no

OpenStudy (anonymous):

Nah trust me, you are gonna be successful

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

its -3.0 FOR THAT NEXT ONE OK

OpenStudy (anonymous):

Wait, so how would I write this?

OpenStudy (anonymous):

ok in f(x)=2.5-10.5 you wirtine in that where 2 is write -5.5 ok

OpenStudy (anonymous):

and g(X)=64(0.58) write 16 ok

OpenStudy (anonymous):

and where 3 is f(x)=2.5-10.5 you witre -3.0 ok

OpenStudy (anonymous):

you write this down

OpenStudy (anonymous):

Sry, i lost connection to openstudy. dont no what happened

OpenStudy (anonymous):

Ok, I am caught up. Now what?

OpenStudy (anonymous):

where g(x)=64(0.5) is wirte 8 on the 3 page ok this would be easy if we were on goolg and video chat i could tell you where to write it down

OpenStudy (anonymous):

Id prefer to do it on here